OpenStudy (anonymous):

How many potassium hydroxide (KOH) formula units are present in 6.89 mol of KOH?:A. 4.1498 × 1024 B. 4.1498 × 1023C. 8.299 × 1024 D. 8.299 × 1023

OpenStudy (anonymous):

PLEASE SOMEONE>>>HELP!!!!!...How many potassium hydroxide (KOH) formula units are present in 6.89 mol of KOH?:A. 4.1498 × 1024 B. 4.1498 × 1023C. 8.299 × 1024 D. 8.299 × 1023

OpenStudy (anonymous):

n=6,89 mol n=N/L --> N= n*L (L is Avogadro's constant) and answer is A